OpenStudy (anonymous):

|(1-2x)/x|=3x. Solve for x?

myininaya (myininaya):

lol i seen this one the other day also

OpenStudy (anonymous):

haha ohh okay

OpenStudy (anonymous):

x = 1/3

myininaya (myininaya):

let me see if can find this one

myininaya (myininaya):

-1 is not a solution

myininaya (myininaya):

just 1/3 is

OpenStudy (anonymous):

x=1/3

OpenStudy (zarkon):

\[|(1-2x)/x|\geq 0\]

myininaya (myininaya):

if x=-1, then 3x is negative

OpenStudy (anonymous):

oh!!there is mod!!!

OpenStudy (anonymous):

check sub x in the equation u'll see that it give the correct answer

myininaya (myininaya):

yes just x=1/3 is correct

OpenStudy (anonymous):

how did you get 3x^2+2x-x-1 when you solved it out?

OpenStudy (anonymous):

who

myininaya (myininaya):

is that 2 suppose to be a 3?

OpenStudy (anonymous):

yea it is

myininaya (myininaya):

we have 3x^2+2x-1=0 i did 3x^2+3x-x-1=0 so i could find the solution(s) by using factor by grouping 3x(x+1)-1(x+1)+0 (x+1)(3x-1)=0 you didn't have to do you could have used the quadratic formula

OpenStudy (anonymous):

ohh okay